OpenStudy (anonymous):

Simplify: Sin(180˚-θ)/sin(90˚-θ) x tan(90˚-θ)

OpenStudy (anonymous):

is this the correct equation of your question? \[\frac{ \sin(180-\theta) }{ \sin(90-\theta)(\tan(90-\theta)) }\]

OpenStudy (anonymous):

if it is, then we can start proving this.

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

is there an x in the denominator?

OpenStudy (anonymous):

Lemme know so I fix the equation.

OpenStudy (anonymous):

oh I see \[\frac{ \sin(180-\theta) }{ \sin(90-\theta) } (\tan(90-\theta))\]

OpenStudy (anonymous):

Sin(180-theta) ? sin(90-theta) multiplied by tan (90-theta)

OpenStudy (anonymous):

like that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, do you know the cofunction identities? \[\sin(90-\theta)= \cos ( \theta )\]

OpenStudy (anonymous):

90 degrees = pi/2 same as \[\sin(\frac{ \pi }{ 2 } - \theta) = \cos(\theta)\]

OpenStudy (anonymous):

and \[\sin(180-\theta) = \sin (\theta)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and \[\tan(90-\theta)= \cot (\theta)\]

OpenStudy (anonymous):

so is that the final answer?

OpenStudy (anonymous):

remember that sin/cos= tan

OpenStudy (anonymous):

oh ok....Thanks heaps

OpenStudy (anonymous):

\(\color{salmon}{\large\heartsuit}\large\text{You're Most Welcome! }\color{salmon}\heartsuit\) \(~~~~~~~~~~~~~~~\color{green}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\) \(~~~~~~~~~~~~~~~~~~~~\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar} \)

OpenStudy (anonymous):

That helps a lot with other questions as well thank you

OpenStudy (anonymous):

hm wait, I think I have a small error

OpenStudy (anonymous):

when I plug in it was supposed to be \[\frac{ \sin(x) }{ \cos(x) } (\cot(x))\] \[=\tan(x) (\frac{ 1 }{ \tan(x) })\] = 1. tan(90-theta) = cot (theta)

OpenStudy (anonymous):

Yes that makes sense........thanks

OpenStudy (anonymous):

:)

OpenStudy (paki):

nice work...

OpenStudy (anonymous):

like how I tagged pantie power instead of you paki, @paki wow..

OpenStudy (anonymous):

:)

OpenStudy (paki):

hahahha yeah :)